OpenStudy (anonymous):

g(x) = 3x + 1 evaluate g(-2)

OpenStudy (texaschic101):

sub in -2 for x g(x) = 3x + 1......g(-2) g(-2) = 3(-2) + 1 g(-2) = -6 + 1 g(-2) = -5